OpenStudy (precal):

need help finding local extrema, inflection points, and x and y intercepts when they exist

OpenStudy (bahrom7893):

post away.

OpenStudy (precal):

\[f(x)=\frac{1}{15}x^3-x+1\]

OpenStudy (precal):

when x is 0 y is 1 so I know that the y-intercept is (0,1)

OpenStudy (bahrom7893):

okay for local extrema: first derivative for inflections: second derivative for y-int: set x = 0, solve for y for x-int: set y=0, solve for x

OpenStudy (precal):

for the first derivative I got \[f \prime (x)=\frac{1}{5}x^2-1\]

OpenStudy (bahrom7893):

Let's find the extrema first: f'(x) = (3/15)x^2-1 = 0 f'(x) = (x^2)/5 - 1 = 0 x^2 - 5 = 0 x^2 = 5 x = +/-5